PLZ tell me with setps


Find value of G in C.G.S using dimesions and units if its value in MKS IS 6.67 *10(raised)-8 dyne.cm(squared)g(raised)-2

Grade:10

1 Answers

Chetan Mandayam Nayakar
12 years ago

in MKS units G=6.67*10-11 Nm2/kg2, one N=kgm/s2 =1000g*100cm/s2 =105 dyne,1m^2=10^4 cm2, 1kg^2=10^6g2,

G=6.67*10^(-11)*105*104/10^6 =6.67*10-8 dyne*cm^2/g^2